Suppose the joint p.d.f of (X,Y) is given by f(x,y) = x^2 + xy/3 0 < x < 1 , 0 < y < 2 = 0 otherwise a) Verify that given f(x,y) is a joint density function? b) Find marginal and conditional probability density functions? c) Compute P (Y < ½ | X < ½ )


1
Expert's answer
2022-01-11T17:35:19-0500

a)

for joint p.d.f:

"\int^{\infin}_{-\infin}\int^{\infin}_{-\infin}f(x,y)dxdy=1"


we have:

"\int^{2}_{0}\int^{1}_{0}(x^2 + xy/3)dxdy=\int^{2}_{0}(x^3/3+x^2y/6)|^1_0dy="


"=\int^{2}_{0}(1/3+y/6)dy=(y/3+y^2/12)|^2_0=2/3+4/12=1"


so, f(x,y) is joint density function


b)

Marginal PDFs:

"f(x)=\int^2_0 f(x,y)dy=\int^2_0 (x^2 + xy/3)dy=(x^2y+xy^2/6)|^2_0="


"=2x^2+2x\/3"


"f(y)=\int^1_0 f(x,y)dx=\int^1_0 (x^2 + xy/3)dx=(x^3/3+x^2y/6)|^1_0="


"=1\/3+y\/6"


conditional probability density functions:

"f(x|y)=\frac{f(x,y)}{f(y)}=\frac{x^2 + xy/3}{1/3+y/6}"


"f(y|x)=\frac{f(x,y)}{f(x)}=\frac{x^2 + xy/3}{2x^2+2x/3}=\frac{x + y/3}{2x+2/3}"


c)

"P (Y < ½ | X < ½ )=\frac{P (Y < ½ ,X < ½ )}{P(X<1/2)}"


"P (Y < ½ ,X < ½ )=\int^{1/2}_{0}\int^{1/2}_{0}f(x,y)dxdy=\int^{1/2}_{0}\int^{1/2}_{0}(x^2 + xy/3)dxdy="


"=\int^{1/2}_{0}(x^3/3+x^2y/6)|^{1/2}_{0}dy=\int^{1/2}_{0}(1/24+y/24)dy="


"=(y\/24+y^2\/48)|^{1\/2}_0=1\/48+1\/192=5\/192"


"P(X<1/2)=\int^{2}_{0}\int^{1/2}_{0}f(x,y)dxdy=\int^{2}_{0}\int^{1/2}_{0}(x^2 + xy/3)dxdy="


"=\int^{2}_{0}(x^3/3+x^2y/6)|^{1/2}_0dy=\int^{2}_{0}(1/24+y/24)dy="


"=(y\/24+y^2\/48)|^2_0=1\/12+1\/12=1\/6"


"P (Y < ½ | X < ½ )=5\cdot6/192=5/32"